OpenStudy (anonymous):

How would I solve y/3 = y+6/8? I know i have to cross multiply so then I get 8y = 3y + 6but what next?

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

my friend helped me actually, i forgot to distribute thanks tho!!

pooja195 (pooja195):

\[\huge~\rm~\bf~\frac{ y }{ 3 }=\frac{ y+6 }{ 8 }\] Cross multiply means what it means :P you actually cross multiply 8y=3(y+6) From here you solve for y

OpenStudy (anonymous):

thank you !!

OpenStudy (jdoe0001):

\(\bf \cfrac{y}{3}=\cfrac{y+6}{8}\implies 8(y)=3(y+6)\implies 8y=3y+18 \\ \quad \\ \textit{then, subtract "3y" from both sides}\)

pooja195 (pooja195):

^I was hoping they would do that part them self so that they would learn. Anyways yw ^_^ Welcome to openstudy!

OpenStudy (anonymous):

yea i understand how to distribute i just forgot to do it

pooja195 (pooja195):

a(b+c)=a*b+a*b 3(y+6)=3*y+3*6